OpenStudy (yacoub1993):

Verify the identity. cot2x + csc2x = 2csc2x - 1

OpenStudy (rajat97):

yeah i'm on it

OpenStudy (rajat97):

csc is cosec right??

OpenStudy (yacoub1993):

i dont know anything about this @rajat97

OpenStudy (rajat97):

okay no problem

OpenStudy (rajat97):

i'm not getting the rhs correct would you please check the identity again??? or if it is to check it , then it is wrong

OpenStudy (rajat97):

i think so

OpenStudy (yacoub1993):

Verify the identity. cot^2x + csc^2x = 2csc^2x - 1

OpenStudy (yacoub1993):

@rajat97 sory my mistake it is ^2 not just 2

OpenStudy (rajat97):

okay now i'm on it and its okay

OpenStudy (rajat97):

okay here csc x is cosecx we have an identity cot^2x +1 =cosec^2x here , lhs=cot^2x+cosec^2x and we can write it as 1+cosec^2x +cosec^2x(from the identity) so cot^2x+cosec^x=1+2cosec^2x

OpenStudy (rajat97):

@Yacoub1993

OpenStudy (yacoub1993):

yes i got it @rajat97

OpenStudy (rajat97):

:)

OpenStudy (rajat97):

any more questions

OpenStudy (yacoub1993):

Verify the identity. 1 + sec^2xsin^2x = sec^2x

OpenStudy (rajat97):

we can write sec^2x=1/cos^2x so sec^2xsin^x=sin^2x/cos^2x=tan^2x and by an identity, 1+tan^2x=sec^2x

OpenStudy (yacoub1993):

how should i write it on the paper

OpenStudy (rajat97):

i'll upload a photo

OpenStudy (yacoub1993):

ok

OpenStudy (rajat97):

wait for a few moments please

OpenStudy (yacoub1993):

ok

OpenStudy (rajat97):

not too clear but i think you can zoom

OpenStudy (yacoub1993):

Verify the identity. cos(α - β) - cos(α + β) = 2 sin α sin β

OpenStudy (yacoub1993):

can you write the same way you did with the last one...

OpenStudy (yacoub1993):

IT IS VERY CLEAR ..

OpenStudy (rajat97):

i'm on it

OpenStudy (yacoub1993):

The following data represents the normal monthly precipitation for a certain city. Draw a scatter diagram of the data for one period. (You do not need to submit the scatter diagram). Find the sinusoidal function of the form that fits the data.

OpenStudy (rajat97):

please wait for a moment i'm on another question (physics) sorry and i'll be back in a few moments

OpenStudy (yacoub1993):

ok

OpenStudy (rajat97):

sorry but i don't have any idea about it any other questions